# An extension of Steinberg’s Theorem to biquotient pairs of subgroups

### v2

@@ 2 Biquotient pairs
In this section, we gather preliminary results on biquotients that facilitate certain reduction steps in the proof of Theorem 3. We begin by recalling the key definitions from the introduction.
@@ 2.1 Reduction to diagonal
$(H_{1}\times H_{2})\backslash\!\backslash G:=(H_{1}\times H_{2})\backslash(G\times G)/\Delta G,$
where $\Delta\colon G\hookrightarrow G\times G$ denotes the diagonal subgroup. Explicitly, the isomorphism is given by $g\mapsto(g,1)$, with inverse $g_{1}g_{2}^{-1}\mapsfrom(g_{1},g_{2})$.
@@
Let $G$ be a compact Lie group with closed subgroups $H_{1}$ and $H_{2}$. The triple $(G,H_{1},H_{2})$ satisfies the lax/strict biquotient condition if and only if the triple $(G\times G,H_{1}\times H_{2},\Delta G)$ does. Irrespective of this condition, $\mathrm{interrank}(H_{1},H_{2})=\mathrm{interrank}(H_{1}\times H_{2},\Delta G)$, and we have canonical isomorphisms of abelian groups
$\operatorname{Tor}_{i}^{\mathrm{R}(G\times G)}(\mathrm{R}(H_{1}\times H_{2}),\mathrm{R}G)\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2}).$
@@
$\operatorname{Tor}_{i}^{\mathrm{R}(G\times G)}(\mathrm{R}(H_{1}\times H_{2}),\mathrm{R}G)_{(\Delta^{*})^{-1}(\mathfrak{p})}\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}},$
in all degrees $i$. Finally,
@@
For the first assertions, note that $(H_{1}\times H_{2})\cap(g_{1},g_{2})\Delta G(g_{1},g_{2})^{-1}$ is isomorphic to $H_{1}\cap g_{1}g_{2}^{-1}H_{2}g_{2}g_{1}^{-1}$ under projection to the first coordinate, and that $\mathrm{Z}(G\times G)=\mathrm{Z}G\times\mathrm{Z}G$.
The isomorphism of Tor-modules can be obtained from Theorem 20 by taking $\Gamma$, $\Lambda$, $\Sigma$, $A$, $B$ and $C$ as indicated in the following diagram:
[…]
The Tor-modules assumed to vanish in positive degrees in Theorem 20 are $\operatorname{Tor}_{*}^{\mathbb{Z}}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ and $\operatorname{Tor}_{*}^{\mathrm{R}G}(\mathrm{R}H_{2},\mathrm{R}G)$. In both cases, this vanishing is clear because at least one of the modules is free over the considered base ring. So we obtain isomorphisms of $\mathrm{R}G\otimes\mathrm{R}G$-modules
@@
in all degrees $i$. To simplify notation, let us temporarily write $G^{2}$ for $G\times G$. Recall that we have a natural isomorphisms $\mathrm{R}(H_{1}\times H_{2})\cong\mathrm{R}H_{1}\otimes\mathrm{R}H_{2}$ and, in particular, $\mathrm{R}(G^{2})\cong\mathrm{R}G\otimes\mathrm{R}G$. We can therefore rewrite the above isomorphisms in the form asserted in the proposition. Under the identification $\mathrm{R}(G^{2})\cong\mathrm{R}G\otimes\mathrm{R}G$, the restriction along $\Delta$ is simply the ring multiplication $\Delta^{*}\colon\mathrm{R}(G^{2})\twoheadrightarrow\mathrm{R}G$. The $\mathrm{R}(G^{2})$-module structure on the right side of the isomorphism is induced from the usual $\mathrm{R}G$-module structure via this epimomorphism $\Delta^{*}$. The second, localized claim follows immediately.
For the final claim, recall again that $\mathrm{Z}(G^{2})=\mathrm{Z}G\times\mathrm{Z}G$. The epimorphism $\Delta^{*}\colon\mathrm{R}(G^{2})\twoheadrightarrow\mathrm{R}G$ restricts to an epimorphism $\mathrm{R}(\mathrm{Z}(G^{2}))\to\mathrm{R}(\mathrm{Z}G)$, which we can view as an epimorphism of $\mathrm{R}(G^{2})$-modules. Localizing at $(\Delta^{*})^{-1}(\mathfrak{p})$ for some prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ yields an epimorphism $\mathrm{R}(\mathrm{Z}(G^{2}))_{(\Delta^{*})^{-1}(\mathfrak{p})}\to(\mathrm{R}(\mathrm{Z}G))_{\mathfrak{p}}$. So if $\mathrm{R}(\mathrm{Z}G)$ is supported at $\mathfrak{p}$, then $\mathrm{R}(\mathrm{Z}(G^{2}))$ must be supported at $(\Delta^{*})^{-1}(\mathfrak{p})$. ∎
@@ 2.2 The biquotient condition can be checked on finitely many intersections
By definition, the biquotient condition is a condition on infinitely many intersections $H_{1}\cap gH_{2}g^{-1}$. We show here that only finitely many intersections need to be taken into account, and that, similarly, the intersection rank of $H_{1}$ and $H_{2}$ can be computed from only finitely many intersections. In short, we can replace $H_{1}$ and $H_{2}$ by tori (Proposition 26), and we need only consider conjugates by elements of the Weyl group of $G$ (Proposition 28).
@@
By definition, $\mathrm{interrank}(H_{1},H_{2})$ is the maximum over the ranks of all tori $S$ contained in the intersections $H_{1}\cap gH_{2}g^{-1}$, where $g$ ranges over the elements of $G$. Any such torus $S$ is conjugate to a subgroup of $T_{1}$ and a subgroup of $T_{2}$, hence conjugate to a subgroup of $T_{1}\cap g^{\prime}T_{2}g^{\prime-1}$ for some $g^{\prime}\in G$. This shows that the interranks agree. If $H_{1}$ and $H_{2}$ satisfy the lax biquotient condition, then clearly so do $T_{1}$ and $T_{2}$. Conversely, suppose $T_{1}$ and $T_{2}$ satisfy the lax biquotient condition, and suppose that $h\in H_{1}\cap gH_{2}g^{-1}$ for some $g\in G$. As any element of $H_{1}$ is conjugate to an element of $T_{1}$, and any element of $H_{2}$ is conjugate to an element of $T_{2}$, we find that $h=h_{1}t_{1}h_{1}^{-1}=gh_{2}t_{2}h_{2}^{-1}g^{-1}$ for certain $h_{i}\in H_{i}$ and $t_{i}\in T_{i}$. It follows that $t_{1}\in T_{1}\cap g^{\prime}T_{2}g^{\prime-1}$ for $g^{\prime}:=h_{1}^{-1}gh_{2}$, hence $t_{1}$ is central by our assumption on $T_{1}$ and $T_{2}$. So $h=h_{1}t_{1}h_{1}^{-1}=t_{1}$ is central. The argument for the strict condition is analogous. ∎
Suppose now that $G$ is compact and connected, and choose a maximal torus $T$ in $G$. The associated Weyl group $\mathrm{W}:=\mathrm{N}_{G}T/T$ is a finite group. Given a subgroup $S\subseteq T$, we write $w(S)$ for the conjugate $hSh^{-1}$ given by an arbitrary representative $h\in\mathrm{N}_{G}T$ of $w\in\mathrm{W}$.
@@
It suffices to show that, for every element $g\in G$, there exists an element $w\in\mathrm{W}$ such that $T\cap gT_{2}g^{-1}\subseteq w(T_{2})$. We show this by adapting arguments from the proof of [BtD:Lie]*IV, Lemma 2.5. Consider $S:=T\cap gT_{2}g^{-1}$. Conjugation by $g$ restricts to an isomorphism between centralizers: $c(g)\colon\mathrm{Z}_{G}(g^{-1}Sg)\to\mathrm{Z}_{G}S$. As $g^{-1}Sg\subseteq T$, and as $T$ is abelian, $T\subseteq\mathrm{Z}_{G}(g^{-1}Sg)$, so $c(g)(T)\subseteq\mathrm{Z}_{G}S$. Likewise, $T\subseteq\mathrm{Z}_{G}S$. So $c(g)(T)$ and $T$ are two maximal tori contained in (the identity component of) $\mathrm{Z}_{G}S$, and hence conjugate in (the identity component of) $\mathrm{Z}_{G}S$. So there exists some element $f\in\mathrm{Z}_{G}S$ such that $fgTg^{-1}f^{-1}=T$, i. e. such that $fg\in\mathrm{N}_{G}T$. Take $w$ to be the image of $fg$ in $\mathrm{W}$. The equality $S=fSf^{-1}$ now expands to $T\cap gT_{2}g^{-1}=f(T\cap gT_{2}g^{-1})f^{-1}$, and the group on the right is clearly contained in $fgT_{2}g^{-1}f^{-1}=w(T_{2})$. ∎
The following proposition is immediate from Lemma 27.
@@ 2.3 Enlarging tori while keeping intersections almost intact
$\operatorname{rank}{G}\geq\operatorname{rank}{H_{1}}+\operatorname{rank}{H_{2}}-\mathrm{interrank}(H_{1},H_{2})$
The following proposition shows that, when $H_{1}$ and $H_{2}$ are tori, we can always enlarge $H_{2}$ such that the above inequality becomes an equality, but without affecting the intersection rank.
@@
Consider a compact connected Lie group $G$ with a maximal torus $T$. Given two subtori $T_{1},T_{2}\subseteq T$, there exists a torus $T_{1}^{+}$ intermediate between $T_{1}$ and $T$ such that
@@
Even when the triple $(G,T_{1},T_{2})$ satisfies the strict biquotient condition, it is not in general possible to find a torus $T_{1}^{+}$ satisfying the above conditions and such that $(G,T_{1}^{+},T_{2})$ still satisfies the biquotient condition. In passing from $T_{1}$ to $T_{1}^{+}$, it can happen that the intersections $T_{1}^{+}\cap w(T_{2})$ acquire additional, non-central points. For example, consider the rank-two group $G:=\mathrm{SU}(3)$ with its standard diagonal torus $T=\{\operatorname{diag}(a,b,a^{-1}b^{-1})\}$, the trivial subtorus $T_{1}:=\{\operatorname{diag}(1,1,1)\}$ and the rank-one subtorus $T_{2}:=\{\operatorname{diag}(a,a^{-1},1)\}$ defined by one of the simple coroots. Clearly, the triple $(G,T_{1},T_{2})$ satisfies the strict biquotient condition. However, any rank-one subtorus $T_{1}^{+}\subseteq T$ intersects at least one of the conjugates $T_{2}$, $T_{2}^{\prime}:=\{\operatorname{diag}(1,a,a^{-1})\}$ and $T_{2}^{\prime\prime}:=\{\operatorname{diag}(a,1,a^{-1})\}$ of $T_{2}$ non-trivially, in a non-central point, so that $(G,T_{1}^{+},T_{2})$ does not even satisfy the lax biquotient condition.
Proof.
@@
Let us abbreviate the rank as follows: $r:=\operatorname{rank}T$, $r_{1}:=\operatorname{rank}T_{1}$, $r_{2}:=\operatorname{rank}T_{2}$, $r^{\prime}:=\mathrm{interrank}(T_{1},T_{2})$. By Proposition 28, $r^{\prime}=\max_{w\in\mathrm{W}}\operatorname{rank}{(T_{1}\cap w(T_{2}))}$. In view of Lemma 27, it suffices to construct an intermediate torus $T_{1}^{+}$ of rank $r_{1}^{+}:=r-r_{2}+r^{\prime}$ such that $\operatorname{rank}{(T_{1}^{+}\cap w(T_{2}))}=r^{\prime}$ for each of the finitely many Weyl group elements $w\in\mathrm{W}$.
@@
The proof works by translating the statement from subgroups of $T$ to quotients of $\mathbb{Z}^{r}$, to subgroups of $\mathbb{Z}^{r}$ and finally to sub-vector spaces of $\mathbb{Q}^{r}$, where it can be solved essentially by linear algebra. The following paragraphs provide some details. To begin with, we observe:
Let $W_{1}$ and $W_{2}$ be sub-vector spaces of $\mathbb{Q}^{r}$, with $c_{i}:=\dim W_{i}$. For any integer $c_{1}^{-}$ such that $0\leq c_{1}^{-}\leq\dim(W_{1}+W_{2})-c_{2}$, there exists a sub-vector space $W_{1}^{-}\subseteq W_{1}$ of dimension $c_{1}^{-}$ such that $\dim(W_{1}^{-}+W_{2})=c_{1}^{-}+c_{2}$. Moreover, the sub-vector spaces $W_{1}^{-}$ with these properties form a (non-empty) Zariski-open subset of rational Grassmannian $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$.
This observation could be phrased more elegantly, and proved, by noting that $\dim(W_{1}+W_{2})-c_{2}=c_{1}-\dim(W_{1}\cap W_{2})$, and by noting that the condition on the dimension of $W_{1}^{-}+W_{2}$ is equivalent to the condition $W_{1}^{-}\cap W_{2}=\{0\}$. Such $W_{1}^{-}$ exists because we can simply choose a complement $W^{\prime}$ of $W_{1}\cap W_{2}$ in $W_{1}$, and then pick an arbitrary subspace $W_{1}^{-}\subseteq W^{\prime}$ of dimension $c_{1}^{-}$. To see that trivial intersection with $W_{2}$ is an open condition on $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$, note that this condition on $W_{1}^{-}$ can be rephrased as: the restriction of the natural projection $W_{1}\twoheadrightarrow W_{1}/(W_{1}\cap W_{2})$ to $W_{1}^{-}$ is a monomorphism. By change of coordinates, we may assume without loss of generality that this projection is the projection onto the first $c_{1}^{-}$ coordinates. Then the restriction is a monomorphism if and only if, for any basis $(b_{1},\dots,b_{c_{1}^{-}})$ of $W_{1}^{-}$, the leading principal $(c_{1}^{-}\times c_{1}^{-})$-minor is non-zero, which is clearly an open condition on the Grassmannian.
Next, we replace the single space $W_{2}$ by finitely many sub-vector spaces $W_{2}^{(i)}$:
Suppose $W_{1}$ and $W_{2}^{(1)},\dots,W_{2}^{(n)}$ are sub-vector spaces of $\mathbb{Q}^{r}$, with all $W_{2}^{(i)}$ of equal dimension. Let $c_{1}:=\dim W_{1}$, $c_{2}:=\dim W_{2}^{(i)}$, $c^{\prime}:=\min_{i}\dim(W_{1}+W_{2}^{(i)})$ and $c_{1}^{-}:=c^{\prime}-c_{2}$. There exists a sub-vector space $W_{1}^{-}\subseteq W_{1}$ such that $c_{1}^{-}=\dim W_{1}^{-}$ and, for each $i$, $c^{\prime}=\dim(W_{1}^{-}+W_{2}^{(i)})$.
Indeed, for each $i$, we have $c_{2}\leq c^{\prime}\leq\dim(W_{1}+W_{2}^{(i)})$, so $0\leq c_{1}^{-}\leq\dim(W_{1}+W_{2}^{(i)})-c_{2}$. Thus, by the previous observation, for each $i$ the sub-vector spaces $W_{1}^{-}\subseteq W_{1}$ with $\dim(W_{1}^{-}+W_{2}^{(i)})=c^{\prime}$ form a non-empty Zariski-open subset of $\mathrm{Gr}_{\mathbb{Q}}(c_{1}^{-},W_{1})$. So we can pick any $W_{1}^{-}$ that lies in the intersection of these finitely many Zariski open subsets.
From the above observation concerning sub-vector spaces of $\mathbb{Q}^{r}$, we deduce an analogous observation concerning direct summands of $\mathbb{Z}^{r}$:
Suppose $K_{1}$ and $K_{2}^{(1)},\dots,K_{2}^{(n)}$ are direct summands of $\mathbb{Z}^{r}$, with all $K_{2}^{(i)}$ of equal rank. Let $c_{1}:=\operatorname{rank}K_{1}$, $c_{2}:=\operatorname{rank}K_{2}^{(i)}$, $c^{\prime}:=\min_{i}\operatorname{rank}(K_{1}+K_{2}^{(i)})$ and $c_{1}^{-}:=c^{\prime}-c_{2}$. There exists a direct summand $K_{1}^{-}$ of $\mathbb{Z}^{r}$ such that $K_{1}^{-}\subseteq K_{1}$ with $c_{1}^{-}=\operatorname{rank}K_{1}^{-}$ and, for each $i$, $c^{\prime}=\operatorname{rank}(K_{1}^{-}+K_{2}^{(i)})$.
Indeed, note that for a sub-vector space $W\subseteq\mathbb{Q}^{r}$, the intersection $W\cap\mathbb{Z}^{r}$ is a direct summand of $\mathbb{Z}^{r}$, of rank equal to the dimension of $W$. Conversely, from any subgroup $K\subseteq\mathbb{Z}^{r}$, we can obtain the sub-vector space $K\otimes_{\mathbb{Z}}\mathbb{Q}\subseteq\mathbb{Q}^{r}$ of dimension equal to the rank of $K$. We obtain in this way a bijection between sub-vector spaces of $\mathbb{Q}^{r}$ and direct summands of $\mathbb{Z}^{r}$. Moreover, for arbitrary subgroups $K_{1},K_{2}\subseteq\mathbb{Z}^{r}$, we have $(K_{1}+K_{2})\otimes\mathbb{Q}=K_{1}\otimes\mathbb{Q}+K_{2}\otimes\mathbb{Q}$. (Note that $K_{1}+K_{2}$ may not be a direct summand even when $K_{1}$ and $K_{2}$ are.) Thus, taking the vector spaces $W_{1}:=K_{1}\otimes\mathbb{Q}$ and $W_{2}^{(i)}:=K_{2}^{(i)}\otimes\mathbb{Q}$ in the previous observation, we obtain a vector space $W_{1}^{-}\subseteq W_{1}$ with the properties specified above, and we may take $K_{1}^{-}:=W_{1}^{-}\cap\mathbb{Z}^{r}$. As noted, this is a direct summand of $\mathbb{Z}^{r}$, clearly contained in $K_{1}$, and the ranks work out as intended. Finally, using Lemma 4, the observation corncerning direct summands of $\mathbb{Z}^{r}$ may be translated into the following statement about tori:

### v3

@@ 2 Biquotient pairs
In this section, we gather preliminary results on biquotients that facilitate certain reduction steps in the proof of Theorem˜3. We begin by recalling the key definitions from the introduction.
@@ 2.1 Reduction to diagonal
[…]
where $\Delta G$ denotes the diagonal subgroup of $G\times G$. Explicitly, the isomorphism is given by $g\mapsto(g,1)$, with inverse $g_{1}g_{2}^{-1}\mapsfrom(g_{1},g_{2})$.
@@
Let $G$ be a compact Lie group with closed subgroups $H_{1}$ and $H_{2}$. Denote as before by $\mathrm{Z}G$ the centre of $G$, and by $\Delta\colon G\hookrightarrow G\times G$ the inclusion of the diagonal subgroup.
  1. (a)
    The triple $(G,H_{1},H_{2})$ satisfies the lax/strict biquotient condition if and only if the triple $(G\times G,H_{1}\times H_{2},\Delta G)$ does.
  2. (b)
    The intersection ranks satisfy the equality
    $\mathrm{interrank}(H_{1},H_{2})=\mathrm{interrank}(H_{1}\times H_{2},\Delta G).$
  3. (c)
    Considering $\mathrm{R}G$ as an $\mathrm{R}(G\times G)$-module via restriction along $\Delta$, we have canonical isomorphisms of abelian groups
    $\operatorname{Tor}_{i}^{\mathrm{R}(G\times G)}\left(\mathrm{R}(H_{1}\times H_{2}),\mathrm{R}G\right)\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$
@@
    $\operatorname{Tor}_{i}^{\mathrm{R}(G\times G)}(\mathrm{R}(H_{1}\times H_{2}),\mathrm{R}G)_{(\Delta^{*})^{-1}(\mathfrak{p})}\cong\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}$
    in all degrees $i$.
  4. (d)
    For a prime ideal $\mathfrak{p}\subset\mathrm{R}G$,
@@
For claims (a) and (b), note that $(H_{1}\times H_{2})\cap(g_{1},g_{2})\Delta G(g_{1},g_{2})^{-1}$ is isomorphic to $H_{1}\cap g_{1}g_{2}^{-1}H_{2}g_{2}g_{1}^{-1}$ under projection to the first coordinate, and that $\mathrm{Z}(G\times G)=\mathrm{Z}G\times\mathrm{Z}G$.
The first isomorphism of Tor-modules in claim (c) can be obtained from Theorem˜21 by taking $\Gamma$, $\Lambda$, $\Sigma$, $A$, $B$ and $C$ as indicated in the following diagram:
$\mathrm{R}H_{1}$$\mathrm{R}H_{2}$$\mathrm{R}G$$\mathbb{Z}$$\mathrm{R}G$$\mathrm{R}G$}
The Tor-modules assumed to vanish in positive degrees in Theorem˜21 are $\operatorname{Tor}_{*}^{\mathbb{Z}}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ and $\operatorname{Tor}_{*}^{\mathrm{R}G}(\mathrm{R}H_{2},\mathrm{R}G)$. In both cases, this vanishing is clear because at least one of the modules is free over the considered base ring. So we obtain isomorphisms of $\mathrm{R}G\otimes\mathrm{R}G$-modules
@@
in all degrees $i$. To simplify notation, let us temporarily write $G^{2}$ for $G\times G$. Recall that we have natural isomorphisms $\mathrm{R}(H_{1}\times H_{2})\cong\mathrm{R}H_{1}\otimes\mathrm{R}H_{2}$ and, in particular, $\mathrm{R}(G^{2})\cong\mathrm{R}G\otimes\mathrm{R}G$. Under the latter isomorphism, the ring multiplication $\mathrm{R}G\otimes\mathrm{R}G\to\mathrm{R}G$ gets identified with the restriction $\Delta^{*}\colon\mathrm{R}(G^{2})\to\mathrm{R}G$. We can therefore rewrite the above isomorphisms in the form asserted in the proposition. The second, localized claim follows immediately.
For claim (d), recall again that $\mathrm{Z}(G^{2})=\mathrm{Z}G\times\mathrm{Z}G$. The surjection $\Delta^{*}\colon\mathrm{R}(G^{2})\twoheadrightarrow\mathrm{R}G$ restricts to a surjection $\mathrm{R}(\mathrm{Z}(G^{2}))\to\mathrm{R}(\mathrm{Z}G)$, which we can view as a surjection of $\mathrm{R}(G^{2})$-modules. Localizing at $(\Delta^{*})^{-1}(\mathfrak{p})$ for some prime ideal $\mathfrak{p}$ of $\mathrm{R}G$ yields a surjection $\mathrm{R}(\mathrm{Z}(G^{2}))_{(\Delta^{*})^{-1}(\mathfrak{p})}\to(\mathrm{R}(\mathrm{Z}G))_{\mathfrak{p}}$. So if $\mathrm{R}(\mathrm{Z}G)$ is supported at $\mathfrak{p}$, then $\mathrm{R}(\mathrm{Z}(G^{2}))$ must be supported at $(\Delta^{*})^{-1}(\mathfrak{p})$. ∎
@@ 2.2 The biquotient condition can be checked on finitely many intersections
By definition, the biquotient condition is a condition on infinitely many intersections $H_{1}\cap gH_{2}g^{-1}$. We show here that only finitely many intersections need to be taken into account, and that, similarly, the intersection rank of $H_{1}$ and $H_{2}$ can be computed from only finitely many intersections. Both arguments proceed as follows: we can replace $H_{1}$ and $H_{2}$ by tori (Proposition˜28), and for tori we need only consider conjugates by elements of the Weyl group of $G$ (Proposition˜30).
@@
By definition, $\mathrm{interrank}(H_{1},H_{2})$ is the maximum over the ranks of all tori $S$ contained in the intersections $H_{1}\cap gH_{2}g^{-1}$, where $g$ ranges over the elements of $G$. Any such torus $S$ is conjugate to a subgroup of $T_{1}$ and a subgroup of $T_{2}$, hence conjugate to a subgroup of $T_{1}\cap g^{\prime}T_{2}g^{\prime-1}$ for some $g^{\prime}\in G$. This shows that the interranks agree. If $H_{1}$ and $H_{2}$ satisfy the lax or strict biquotient condition, then clearly so do $T_{1}$ and $T_{2}$. Conversely, suppose $T_{1}$ and $T_{2}$ satisfy the lax (or strict) biquotient condition, and suppose that $h\in H_{1}\cap gH_{2}g^{-1}$ for some $g\in G$. As any element of $H_{1}$ is conjugate to an element of $T_{1}$, and any element of $H_{2}$ is conjugate to an element of $T_{2}$, we find that $h=h_{1}t_{1}h_{1}^{-1}=gh_{2}t_{2}h_{2}^{-1}g^{-1}$ for certain $h_{i}\in H_{i}$ and $t_{i}\in T_{i}$. It follows that $t_{1}\in T_{1}\cap g^{\prime}T_{2}g^{\prime-1}$ for $g^{\prime}:=h_{1}^{-1}gh_{2}$, hence $t_{1}$ is central (or trivial) by our assumption on $T_{1}$ and $T_{2}$. So $h=h_{1}t_{1}h_{1}^{-1}=t_{1}$ is central (or trivial, respectively). ∎
Suppose now that $G$ is compact and connected, with a chosen maximal torus $T$ and associated Weyl group $\mathrm{W}=(\mathrm{N}_{G}T)/T$ as in section˜1.3. Recall that $\mathrm{W}$ is finite. Given a subgroup $S\subseteq T$, we write $w(S)$ for the conjugate $hSh^{-1}$ given by an arbitrary representative $h\in\mathrm{N}_{G}T$ of $w\in\mathrm{W}$.
@@
It suffices to show that, for every element $g\in G$, there exists an element $w\in\mathrm{W}$ such that $T\cap gT_{2}g^{-1}\subseteq w(T_{2})$. We show this by adapting arguments from the proof of [BtD:Lie]*IV, Lemma 2.5. Consider $S:=T\cap gT_{2}g^{-1}$. Conjugation by $g$ restricts to an isomorphism between centralizers: $c(g)\colon\mathrm{Z}_{G}(g^{-1}Sg)\to\mathrm{Z}_{G}S$. As $g^{-1}Sg\subseteq T$, and as $T$ is abelian, $T\subseteq\mathrm{Z}_{G}(g^{-1}Sg)$, so $c(g)(T)\subseteq\mathrm{Z}_{G}S$. Likewise, $T\subseteq\mathrm{Z}_{G}S$. So $c(g)(T)$ and $T$ are two maximal tori contained in (the identity component of) $\mathrm{Z}_{G}S$, and hence conjugate in $\mathrm{Z}_{G}S$. So there exists some element $f\in\mathrm{Z}_{G}S$ such that $fgTg^{-1}f^{-1}=T$, i. e., such that $fg\in\mathrm{N}_{G}T$. Take $w$ to be the image of $fg$ in $\mathrm{W}$. The equality $S=fSf^{-1}$ now expands to $T\cap gT_{2}g^{-1}=f(T\cap gT_{2}g^{-1})f^{-1}$, and the group on the right is clearly contained in $fgT_{2}g^{-1}f^{-1}=w(T_{2})$. ∎
The following proposition is immediate from Lemma˜29.
@@ 2.3 Enlarging tori while keeping intersections almost intact
$\operatorname{rank}{G}\geq\operatorname{rank}{H_{1}}+\operatorname{rank}{H_{2}}-\mathrm{interrank}(H_{1},H_{2}).$
The following proposition shows that, when $H_{1}$ and $H_{2}$ are tori, we can always enlarge $H_{1}$ in such a way that the above inequality becomes an equality, but without affecting the intersection rank.
@@
Consider a compact connected Lie group $G$. Given two tori $T_{1},T_{2}\subseteq G$, there exists a torus $T_{1}^{+}$ containing $T_{1}$ such that
@@
The proof will show that, given a maximal torus $T$ of $G$ containing $T_{1}$, we can choose $T_{1}^{+}$ to be intermediate between $T_{1}$ and $T$, i. e., such that $T_{1}\subseteq T_{1}^{+}\subseteq T$.
Remark 33.
@@
Even when the triple $(G,T_{1},T_{2})$ satisfies the strict biquotient condition, it is not in general possible to find a torus $T_{1}^{+}$ satisfying the above conditions and such that $(G,T_{1}^{+},T_{2})$ still satisfies one of the biquotient conditions. In passing from $T_{1}$ to $T_{1}^{+}$, it can happen that the intersections $T_{1}^{+}\cap w(T_{2})$ acquire additional, non-central points. For example, consider the rank-two group $G:=\mathrm{SU}(3)$ with its standard diagonal torus $T=\{\operatorname{diag}(a,b,a^{-1}b^{-1})\}$, the trivial subtorus $T_{1}:=\{\operatorname{diag}(1,1,1)\}$ and the rank-one subtorus $T_{2}:=\{\operatorname{diag}(a,a^{-1},1)\}$ defined by one of the simple coroots. Clearly, the triple $(G,T_{1},T_{2})$ satisfies the strict biquotient condition. However, any rank-one subtorus $T_{1}^{+}\subseteq T$ intersects at least one of the conjugates $T_{2}$, $T_{2}^{\prime}:=\{\operatorname{diag}(1,a,a^{-1})\}$ and $T_{2}^{\prime\prime}:=\{\operatorname{diag}(a,1,a^{-1})\}$ of $T_{2}$ non-trivially, in a non-central point, so that $(G,T_{1}^{+},T_{2})$ does not even satisfy the lax biquotient condition.
Proof of Proposition˜31.
Choose a maximal torus $T$ of $G$ such that $T_{1}\subseteq T$. Replacing $T_{2}$ by a conjugate subgroup if necessary, we may assume that also $T_{2}\subseteq T$. Then by Proposition˜30,
$\mathrm{interrank}(T_{1},T_{2})=\max_{w\in\mathrm{W}}\operatorname{rank}{(T_{1}\cap w(T_{2}))}.$
@@
Let us abbreviate the ranks as follows: $r:=\operatorname{rank}T$, $r_{1}:=\operatorname{rank}T_{1}$, $r_{2}:=\operatorname{rank}T_{2}$, $r^{\prime}:=\mathrm{interrank}(T_{1},T_{2})$. In view of Lemma˜29, it suffices to construct a torus $T_{1}^{+}$ intermediate between $T_{1}$ and $T$ and of rank $r_{1}^{+}:=r-r_{2}+r^{\prime}$ such that $\operatorname{rank}{(T_{1}^{+}\cap w(T_{2}))}=r^{\prime}$ for each of the finitely many Weyl group elements $w\in\mathrm{W}$.
@@
The proof works by translating the statement from subgroups of $T$ to subgroups of $\mathbb{Z}^{r}$ to vector subspaces of $\mathbb{Q}^{r}$, where it can be solved essentially by linear algebra. The following paragraphs provide some details.
First, let us abstract away the Weyl group and formulate our claim more generally as follows.